# ParisLuco3D: A high-quality target dataset for domain generalization of LiDAR perception

### v2

@@ I Introduction
While open-source datasets display a large variety of scenes and sensor setups, crucial discrepancies have been identified among label sets [3, 4]. Due to this label difference, it is not possible to leverage current online benchmarks to ensure fair comparisons across methods.
@@
SemanticPOSS [8] 2020 N/A ${}^{\dagger}$$\times$ $\times$
@@
ParisLuco3D (Ours) 2023 N/A ${}^{\dagger}$ France urban 1.33 0.16 32
TABLE I: Summary of the various existing LiDAR perception datasets. ${}^{\dagger}$size of the full dataset, as there is no designated split. KITTI provides LOD annotation only for a subset of the point cloud. KITTI-360 provides the LSS annotation on the accumulated and subsampled point cloud rather than on each scan (LSS = LiDAR Semantic Segmentation and LOD = LiDAR Object Detection).
@@
  • The release of a novel dataset, in addition to online benchmarks, to compute cross-domain performances fairly.
@@ II-A1 Datasets for LiDAR semantic segmentation
Thereafter, SemanticKITTI was expanded into KITTI-360 [13], which includes data on the same suburb but on a larger scale.
Subsequently, a few object detection datasets were expanded to incorporate semantic segmentation annotation, such as nuScenes [7] and Waymo [9]. They were acquired in cities and suburbs in the US and Asia. Contrary to SemanticKITTI, nuScenes focuses its annotations on vehicles with a limited number of background classes.
@@ III-D Labels for object detection
ParisLuco3D ${}^{**}$ 64k 14.7 6.4 56.3
TABLE III: Comparison between 3D object detection datasets. All values are computed for the validation splits of the datasets. $*{}^{*}$he 23 original classes of nuScenes are grouped during evaluation into 10 due to a strong similarity in a few fine classes—for example, standing pedestrians and sitting pedestrians. ${}^{**}$hile we annotate 11 different classes, we evaluate 7 of these classes for models trained on nuScenes and 5 for those trained on ONCE, using the intersection of classes between these datasets and ours.
@@ IV-A2 Experiments
[…]
@@ IV-A3 Results and analysis
[…]
@@ IV-B1 Related work
The best performing methods typically use two-steps processes of initial proposals and refinement using features interpolated from keypoints [35, 2]. Recently, methods have been exploring predictions based on object centers [2] as well as transformer-based architectures [36].
@@ IV-B2 Experiments
We evaluated the generalization ability of five different 3D object detection models across the various choices of source datasets: SECOND [32], PointRCNN [34], PointPillars [33], PV-RCNN [35], and CenterPoint [2]. These models were selected due to their availability and their performance across different 3D object detection benchmarks.
@@ IV-B3 Results & analysis
We present our benchmark results in Table VI in terms of mAP. Our models were trained without the intensity channel, as we emphasize in Table VII how intensity hinders the generalization ability of models. In Table VI, despite its relatively low source mAP, the PointPillars models tends to generalize better on ParisLuco3D when trained on ONCE. We believe that this is because it contains a lower number of parameters, and so it tends to be less susceptible to overfitting. On the other hand, PV-RCNN works best when training using nuScenes, as the model is natively capable of capturing more fine-grained details, which is necessary when working on the low resolution of nuScenes.
Table VI highlights the fact both RayDrop and IBN-Net are insufficient generalization baselines and their impact may depend on the source dataset used. For example, using the RayDrop augmentation for a PointPillars model trained on ONCE when the source LiDAR density is higher than that of ParisLuco3D tends to help the generalization ability of this model ($+0.9$ mAP). However, the opposite occurs when we use RayDrop with nuScenes ($-1.2$ mAP). RayDrop increase the generalization for most classes for our PointPillars model trained on ONCE, making it the model with the highest mAP on ParisLuco3D despite having one of the lowest mAP performances on its own source datasets. We believe this highlights the fact that aligning the LiDAR resolutions while taking advantage of large datasets can be a strong generalization method. However, as shown using nuScenes as a training set, it is not a one size fits all method and is dependant on the difference in LiDAR sensor resolutions between a source and a target dataset, which is normally unknown in the conditions of the domain generalization task.
@@
[…]
SECOND [32]
@@
In Table VII, we trained SECOND, CenterPoint, and PointPillars on the ONCE and nuScenes datasets both with and without intensity. It is evident that unlike in semantic segmentation, using intensity for object detection when training on nuScenes actually hinders the final performance by a large amount, ranging from a +0.9 to -10.2 mAP difference on ParisLuco3D compared to models trained without intensity (yet nuScenes and ParisLuco3D have the same LiDAR sensor). We believe this is due to a difference in the intensity distributions of objects, which has been shown to be a major factor of model overfitting for LOD [37].

### v3

@@ I Introduction
While open-source datasets display a large variety of scenes and sensor setups, crucial discrepancies have been identified among label sets [3, 4]. Moreover, despite previous works managing to provide meaningful insights in the domain generalization task, the lack of consensual labels sets can result in unfair comparisons.
@@
SemanticPOSS [8] 2020 N/A 2988 $\times$ $\times$
@@
ParisLuco3D (Ours) 2023 N/A 7501 France urban 1.33 0.16 32
TABLE I: Summary of the various existing LiDAR perception datasets. size of the full dataset, as there is no designated split. KITTI provides LOD annotation only for a subset of the point cloud. KITTI-360 provides the LSS annotation on the accumulated and subsampled point cloud rather than on each scan (LSS = LiDAR Semantic Segmentation and LOD = LiDAR Object Detection).
@@
  • The release of a novel dataset, in addition to online benchmarks, to compute cross-domain performances fairly. This dataset was acquired in the center of Paris, a relatively unavailable scene type.
@@ II-A1 Datasets for LiDAR semantic segmentation
Thereafter, SemanticKITTI was expanded into KITTI-360 [13], which includes data on the same suburb but on a larger scale. Subsequently, a few object detection datasets were expanded to incorporate semantic segmentation annotation, such as nuScenes [7] and Waymo [9]. They were acquired in cities and suburbs in the US and Asia. Contrary to SemanticKITTI, nuScenes focuses its annotations on vehicles with a limited number of background classes.
@@ III-D Labels for object detection
ParisLuco3D 11/7/5∗∗ 64k 14.7 6.4 56.3
TABLE III: Comparison between 3D object detection datasets. All values are computed for the validation splits of the datasets. The 23 original classes of nuScenes are grouped during evaluation into 10 due to a strong similarity in a few fine classes—for example, standing pedestrians and sitting pedestrians. ∗∗While we annotate 11 different classes, we evaluate 7 of these classes for models trained on nuScenes and 5 for those trained on ONCE, using the intersection of classes between these datasets and ours.
@@ IV-A2 Experiments
Model Input type mIoUSK mIoUPL mIoUNS mIoUPL
@@ IV-A3 Results and analysis
For nuScenes to ParisLuco3D, the results were computed with the intensity channel, as the same LiDAR sensor is used in both datasets. This is also illustrated in Table V, where we can see that the results are better with than without intensity. The low mIoU (39.7 mIoUPL for the best model) comes from the trailer and traffic-cone categories for which there are very few examples in ParisLuco3D. It is important to keep these minority classes because they allow models to be evaluated on the rare edge cases.
@@ IV-B1 Related work
The best performing methods typically use two-steps processes of initial proposals and refinement using features interpolated from keypoints [35, 2]. Recently, methods have been exploring predictions based on object centers [2], sparse features [36] as well as transformer-based architectures [37].
@@ IV-B2 Experiments
We evaluated the generalization ability of five different 3D object detection models across the various choices of source datasets: SECOND [32], PointRCNN [34], PointPillars [33], PV-RCNN [35], CenterPoint [2], VoxelNeXt [36], and DSVT [37]. These models were selected due to their availability and their performance across different 3D object detection benchmarks. Notably, VoxelNeXt and DSVT are the current state-of-the-art 3D object detection models for the Argoverse2 [38] dataset and nuScenes [7] dataset respectively.
@@ IV-B3 Results & analysis
We present our benchmark results in Table VI in terms of mAP. Our models were trained without the intensity channel, as we emphasize in Table VII how intensity hinders the generalization ability of models. In Table VI, despite its relatively low source mAP, the PointPillars models tends to generalize better on ParisLuco3D when trained on ONCE. We believe that this is because it contains a lower number of parameters, and so it tends to be less susceptible to overfitting. On the other hand, PV-RCNN works best when training using nuScenes, as the model is natively capable of capturing more fine-grained details, which is necessary when working on the low resolution of nuScenes. Finally, we find that though DSVT achieves high source accuracy on nuScenes when using close to optimal hyperparameters, it suffers the most from generalization accuracy degradation on our dataset.
Table VI highlights the fact both RayDrop and IBN-Net are insufficient generalization baselines and their impact may depend on the source dataset used. For example, using the RayDrop augmentation for a PointPillars model trained on ONCE when the source LiDAR density is higher than that of ParisLuco3D tends to help the generalization ability of this model ($+0.9$ mAP). However, the opposite occurs when we use RayDrop with nuScenes ($-1.4$ mAP). RayDrop increases the generalization for most classes for our PointPillars model trained on ONCE, making it the model with the second highest mAP on ParisLuco3D despite having one of the lowest mAP performances on its own source datasets. We believe this highlights the fact that aligning the LiDAR resolutions while taking advantage of large datasets can be a strong generalization method. However, as shown using nuScenes as a training set, it is not a one size fits all method and is dependant on the difference in LiDAR sensor resolutions between a source and a target dataset, which is normally unknown in the conditions of the domain generalization task.
@@
Model mAPON mAPPL mAPNS mAPPL
PointPillars [33]
@@
In Table VII, we trained PointPillars, CenterPoint, and DSVT on the ONCE and nuScenes datasets both with and without intensity. It is evident that unlike in semantic segmentation, using intensity for object detection when training on nuScenes actually hinders the final performance by a large amount, ranging from a -2.0 to -10.2 mAP difference on ParisLuco3D compared to models trained without intensity (yet nuScenes and ParisLuco3D have the same LiDAR sensor). We believe this is due to a difference in the intensity distributions of objects, which has been shown to be a major factor of model overfitting for LOD [39].